OpenStudy (anonymous):

What is the solution set of |2x – 3| – 6 = –1?

OpenStudy (anonymous):

You first add 6 to the -1

OpenStudy (hba):

did u try to solve it

Parth (parthkohli):

Start by adding 6 to both sides. \( \color{Black}{\Rightarrow |2x - 3| = 5}\) Now put that plus/minus and remove the absolute value. \( \color{Black}{\Rightarrow 2x - 3 = \pm 5}\)

OpenStudy (anonymous):

2x-3-6=-1 2x-9=-1 2x=8 x=4 -2x+3-6=-1 -2x+3=-1 -2x=-4 x=2

Parth (parthkohli):

Solve the equations: \(2x - 3 = +5\) \(2x - 3 = -5\)

Parth (parthkohli):

@sai1234 You skipped the absolute value right there.